# Automatic Derivation of an Optimal Task Frame for Learning and Controlling Contact-Rich Tasks

### v1

@@ Abstract
This study investigates learning from demonstration (LfD) for contact-rich tasks. The procedure for choosing a task frame to express the learned signals for the motion and interaction wrench is often omitted or using expert insight. This article presents a procedure to derive the optimal task frame from motion and wrench data recorded during the demonstration. The procedure is based on two principles that are hypothesized to underpin the control configuration targeted by an expert, and assumes task frame origins and orientations that are fixed to either the world or the robot tool. It is rooted in screw theory, is entirely probabilistic and does not involve any hyperparameters. The procedure was validated by demonstrating several tasks, including surface following and manipulation of articulated objects, showing good agreement between the obtained and the assumed expert task frames. To validate the performance of the learned tasks by a UR10e robot, a constraint-based controller was designed based on the derived task frames and the learned data expressed therein. These experiments showed the effectiveness and versatility of the proposed approach. The task frame derivation approach fills a gap in the state of the art of LfD, bringing LfD for contact-rich tasks closer to practical application.
@@
Learning from demonstration, contact-rich task, task frame, screw theory, constraint-based control
@@ I Introduction
Contact-rich tasks in robotics are object manipulation tasks that involve contact with the environment or with other objects. In this paper, we focus on the manipulation of a single object attached to the robot end effector in contact with another object attached to the environment. Such tasks are abundant in all fields of robotics: industry, domestic robots, health care, etc. They include, for example, the assembly of parts, machining operations during production, pick-and-place actions, and the opening, closing, or handling of objects. Programming contact tasks on robot manipulators can be challenging since it requires coordinating and controlling both the desired motion of an object and its desired interaction force with the environment.
The design of a robot controller for a contact task can be greatly simplified if a suitable reference frame for the control actions of the robot can be found. We refer to such a frame as a task frame. The task frame is defined as a coordinate frame, with an origin and an orientation, in which the motion and force signals are controlled in different directions towards desired values [1]. Given a suitable choice for the task frame, some of these signals become constant or zero, facilitating the design of a stable and robust controller.
@@ I-A Problem and Research Objective
The choice of where to place the task frame depends on an expert’s intuition for the application at hand. Sometimes, the frame is best defined with respect to the robot’s tool. Other times, it is better defined with respect to an object in the world. Considering the orientation, sometimes the axes of the frame are best chosen in the main direction of the translation or the force, and other times they should be in the directions of rotation or moment. To the best of the authors’ knowledge, the underlying principles for choosing a task frame have nowhere been made explicit. This choice is not trivial since it requires insight in the robotic application.
Learning from Demonstration (LfD) offers an attractive alternative to robot programming since it enables non-experts to teach a robot how to do a task without deep knowledge of robotics. The main idea in LfD is to record human demonstrations of a task and build a model from the demonstration data. This model can then be used by the robot to perform the task and adapt to new circumstances.
The LfD approach has not only been applied to pure motion tasks, but to contact tasks as well. Approaches have been proposed to specify the controller by identifying the main directions of force and motion. These directions can be considered for the orientation of the task frame. However, often only the position and force are considered, neglecting the orientation of the object and the moment. Furthermore, existing approaches are often limited in the number of degrees of freedom. Usually only articulated objects are considered with one degree of freedom in motion, such as hinges or sliders.
The objective of this paper is to uncover and establish the underlying principles with which experts choose the task frame. Based on these principles, the aim is to automatically derive a suitable task frame in a data-driven way from recorded motion and force signals instead of relying on an expert’s intuition.
@@ I-B Approach
To derive an optimal task frame, we first establish two principles for an optimal choice of the task frame based on decoupling. Our assumptions are that the data are segmented and that the task frame is either fixed to the world or the tool. Based on these principles we identify multiple candidate origins and orientations for the task frame from the demonstrated motion and wrench data. These candidates are all combined until we arrive at the optimal task frame origin and orientation. Once the task frame has been derived, the desired motion and wrench signals are expressed in this frame and a controller can be defined to perform the task. The automatic task frame derivation procedure is validated for a diverse set of contact-rich tasks, without any assumption on the type of task, by comparing the result with an expert’s choice. The tasks are performed on the robot after designing a constraint-based controller based on the derived task frame. The controller’s robustness and generalization capabilities are tested under disturbances, parameter variations, and alterations in the task frame.
@@ I-C Contribution and Outline
The main contribution of this research is a multi-step algorithm to derive an optimal task frame from demonstrated motion and wrench data. The main discerning features of our task frame derivation approach are that it incorporates data not only of the object’s position and the interaction force, but also of the object’s orientation and the interaction moment. It decides the most appropriate reference frame separately for the origin and orientation of the task. Finally, it decides the most appropriate variable in which the motion and force signals should be parameterized. The proposed method is intended as a tool by which a non-expert user can program contact tasks by first demonstrating the task. This helps to reduce the programming cost and effort of contact tasks, making it more widely accessible.
The remainder of this article is organized as follows. Section II reviews the literature on learning of contact tasks. Section III presents mathematical preliminaries, while Section IV proposes the methodology to automatically derive the optimal task frame from motion and force data. Sections V to VI cover the controller specification and the experimental validation on a variety of contact tasks. Finally, Section VII outlines the major contributions, limitations, and conclusions of the work.
II Related work
@@ II-A Collection of Demonstration Data
Four common ways to collect demonstration data for LfD methods are kinesthetic teaching [2, 3, 4, 5, 6, 7, 8], teleoperation [9, 10], passive observation [11, 12] and simulation. The first method involves physically grasping the robot and moving it through the task, recording motion via the robot’s encoders and interaction wrench via torque sensors at the joints or via a force/torque sensor at the wrist. While cost-effective and fast, this method can be challenging for heavy robots [10], bulky robots [13], redundant robots, dual-arm setups [14], or non-manipulator robots [15], as undesirable dynamics induced by the operator may affect the recorded data [11]. The second method is usually more expensive since it involves an external interface between the robot and the operator, such as a joystick, haptic device or graphical user interface. Collecting demonstration data by teleoperation is also more challenging than kinesthetic teaching [16]. In the third method, the operator uses an instrumented tool to perform the task and record motion and wrench data. In the last method, simulation environments like MoJuCo [17] and Bullet [18] are used to generate demonstration data. The latter method offers abundant data which can enable reinforcement-learning-based methods to widely explore the task without any risk for the robot. However, transferring learned skills from the simulator to the real world is still challenging [19]. Consequently, building a reliable model derived from one [2, 3, 20] or a few [12, 4, 7] demonstrations is a critical aspect of learning contact-rich tasks. In this article, passive observation is employed to record demonstrations as it aligns with the most natural way for the operator to perform tasks [13], but the developed approach is not limited to this demonstration method.
@@ II-B Assumptions and Applications
Some researchers included methods for segmentation in their proposed pipelines for processing the demonstration data, such as [7, 11, 2, 20], but we left segmentation out of the scope of the present paper. I.e., we assume demonstration data has already been segmented and propose an approach to build a task model and controller for a single task segment in which the contact situation remains the same over time. On the other hand, while a lot of approaches in literature focus on LfD for a single task or a limited class of tasks, our aim was to develop a generic approach, without further knowledge or assumptions about the task other than that the contact is between only two objects with a measurable relative motion.
@@ II-C Selection of Task Frame
Task frame selection regained interest in the context of LfD, since demonstration data and task models have to be represented in an appropriate reference frame. There are two common ways to choose a task frame. The first way is, still as in earlier works, to rely on implicit high-level robotics expertise, i.e., without using any prescribed method. The second way is to select between several task frame candidates that are predefined based on expert knowledge. This selection is based on criteria like consistency [7], variance [26] or generalization capability of the demonstration data [27]. A first set of such studies directly select between different task frame candidates [7, 26, 28] or optimize the position and orientation of such candidates [8]. Other studies indirectly learn an appropriate reference frame or components thereof. For example, Manschitz et al. [27] proposed a method which learns to activate a set of attractors that are each defined in a different appropriate reference frame, while Suomalainen and Kyrki [29] identified the motion direction toward the target and one or two compliant axes from demonstrated position and force trajectories, which are then used to design an impedance controller.
In this paper we propose a generic approach to derive an optimal task frame that applies to contact-rich tasks (i.e., not only to motions in free space), does not use any application-specific expert knowledge nor any hyperparameter. The origin and orientation of the task frame can be fixed in different viewpoints (world or tool) so that the task frame as a whole is not necessarily fixed to either world or tool. We use both motion and wrench data, consider both their translational and rotational vector components and automatically derive the motion and wrench vector components of interest from the demonstration data. While all works mentioned above exhibit some of these features, they lack several others. On the other hand, some works deal with aspects that are not the focus of this paper. For example, in [7], at each time step, the proposed method can decide between two different predefined reference frames to express the signals. Furthermore, although it is not common to consider a moving task frame for contact-rich tasks in the context of LfD, this was done in [5].
As final remarks, using a task frame is the most common approach in the literature for LfD, with several researchers proposing modifications or extensions of the task frame concept; however, their work is limited to tasks involving motion in free space and therefore not considered in this overview. Some works use a completely different approach, i.e., other than task frame based. For example, Vochten et al. [30] introduced two types of shape descriptors to represent the motion of the object of interest, along with invariant reference frames, and in a subsequent work [31] extended this idea to the wrench domain. However, each domain still has its own reference frame, and they are not merged as we do in this paper.
@@ II-D Task Model and Control
The task model proposed in this paper is derived from 6-D motion and 6-D wrench data. It consists of 1) a task frame, with its location and orientation as parameters, and 2) the demonstration data, expressed in the task frame and averaged to find the desired values for the motion and interaction wrench. No further processing of the data is performed. A generic constraint-based controller is proposed to follow these desired values. However, we believe that other learning methods could leverage this automatically derived task frame, which is obtained in a data-driven way, for building their models and exploring alternative controllers.
@@ III-A Orientation and Pose
Figure 1 visualizes the similarity transformation under general conditions with three moving frames.
@@ III-B Velocity kinematics
The velocity kinematics of an orientation frame are characterized by the rotational velocity vector $\boldsymbol{\omega}$. When the vector’s coordinates are expressed in $\left[a\right]$, the relation between ${}_{a}\boldsymbol{\omega}$ and the derivative of the rotation matrix $\prescript{b}{}{\boldsymbol{R}}_{a}^{\prime}$ is:
$\prescript{b}{}{\boldsymbol{R}}_{a}^{\prime}=[{}_{a}\boldsymbol{\omega}]_{\times}\prescript{b}{}{\boldsymbol{R}}_{a}.$ (10)
The velocity kinematics of a pose frame are characterized by the twist $\boldsymbol{\mathcalboondox{t}}=(\boldsymbol{\omega}^{T}~{}~{}\boldsymbol{v}^{o\,T})^{T}\in\mathbb{R}^{6\times 1}$, consisting of the rotational velocity vector $\boldsymbol{\omega}$ and the translational velocity vector $\boldsymbol{v}^{o}$ at the origin $o$ of the reference frame. When the twist coordinates are expressed in $\{a\}$, the relation between ${}_{a}\boldsymbol{\mathcalboondox{t}}$ and the time-derivative of the pose matrix $\prescript{b}{}{\boldsymbol{T}}_{a}^{\prime}$ can be worked out as:
$\prescript{b}{}{\boldsymbol{T}}_{a}^{\prime}=[{}_{a}\boldsymbol{\mathcalboondox{t}}]_{\times}\prescript{b}{}{\boldsymbol{T}}_{a}.$ (11)
@@ III-D Time-invariant motion and wrench trajectories
These trajectories can be made time-invariant by re-parameterizing them as a function of a geometric progress variable $\xi$. There are multiple valid choices for the progress variable. For example, it can be defined as the cumulative arc length of the trajectory traversed by a chosen reference point, or as the cumulative rotation angle, or by some combination of the two. Often, the progress variable is also made non-dimensional (ranging from 00 to $1$) so that trajectories of different length can be compared.
Once a progress $\xi(t)$ has been defined, the reparameterizations of pose, wrench, and twist are defined as follows [31]:
$\prescript{}{}{\boldsymbol{T}}{}(\xi)=\prescript{}{}{\boldsymbol{T}}{}(t(\xi)),~\boldsymbol{\mathcalboondox{w}}(\xi)={\boldsymbol{\mathcalboondox{w}}}(t(\xi)),~{}\boldsymbol{\mathcalboondox{t}}(\xi)=\frac{\boldsymbol{\mathcalboondox{t}}(t(\xi))}{\dot{\xi}(t(\xi))}.$ (15)
In practice, for discrete-time data, these reparameterizations are performed using numerical interpolation.
Throughout this paper and in order not to overload notation, we do not explicitly mention the dependence of signals on time or on another progress variable when such dependence is obvious from the context.

### v3

@@ Abstract
In previous work on learning and controlling contact-rich tasks, the procedure for choosing a proper reference frame to express learned signals for the motion and the interaction wrench is often implicit, requires expert insight, or starts from proposed frame candidates. This article presents an automatic method to derive the optimal reference frame, referred to as optimal task frame, directly from the recorded motion and wrench data of the demonstration. Using screw theory, several origin and orientation candidates are generated that maximize decoupling in the data. These candidates are then processed probabilistically, without needing hyperparameters, to obtain the optimal task frame. Its origin and orientation are independently fixed to either the world or the robot tool. The method works regardless of whether the task involves translation, rotation, force, or moment, or any combination thereof. The method was validated for various tasks, including surface following and manipulation of articulated objects, showing good agreement between derived and assumed expert task frames. To validate the robot’s performance, a constraint-based controller was designed based on the data expressed in the derived task frames. These experiments demonstrated the approach’s effectiveness and versatility. The automatic task frame derivation approach supports learning methods to design controllers for a wide range of contact-rich tasks.
@@
Learning from demonstration, contact-rich task, task frame, screw theory, constraint-based control.
@@ I Introduction
Contact-rich tasks in robotics are object manipulation tasks that involve contact with the environment or with other objects. In this paper, we focus on the manipulation of a single object attached to the robot end effector in contact with another object attached to the environment. Such tasks are abundant in all fields of robotics: industry, domestic robots, health care, etc. They include, for example, assembly, machining, picking and placing, and handling of articulated objects such as doors or drawers.
Programming contact tasks on robot manipulators can be challenging since it requires coordinating and controlling both the desired motion of an object that is manipulated and its desired interaction wrench with the environment. Depending on the task, the translational and/or the rotational motion may be relevant, as well as the interaction force and/or the interaction moment.
Learning from Demonstration (LfD) can help to make the programming of contact tasks more intuitive since it allows non-expert users to teach tasks to robots without a deep knowledge of robotics [1]. Fig. 1a illustrates a typical LfD workflow consisting of first recording the data of human demonstrations of the task, then learning a task model from the data and designing a corresponding controller, and finally using this controller on the robot to perform the task while adapting to new circumstances.
Figure 1: (a) The typical workflow in robot task learning consists of data recording, learning a task model and controller from the data, and performing the task on a robot. (b) Our approach adds the automatic derivation of an optimal task frame from the data, prior to learning the task model.
For object manipulation tasks in contact, it is important to define the task controller in a suitable reference frame. We refer to such a frame as a task frame. The task frame is defined as a coordinate frame, with an origin and an orientation, in which the motion and wrench signals are controlled in different directions towards desired values [2]. Given a suitable choice for the task frame, some of these signals become constant or zero, facilitating the design of a stable and robust controller. Therefore, the choice of the task frame is crucial for the success of the control strategy.
@@ I-A Problem and Research Objective
It is not always clear which choice of task frame is best for the controller. In some learning methods, the task frame is chosen implicitly to be the same as the reference frame in which the data were recorded. However, this usually does not lead to an optimal choice. Alternatively, the task frame can be chosen based on the expert user’s intuition for the application at hand. For some tasks, the frame is best defined with respect to the robot’s tool. Other times, it is better defined with respect to the environment. Considering the orientation, sometimes the axes of the frame are best chosen in the main direction of the translation or the interaction force, and other times they should be in the directions of rotation or interaction moment.
Some LfD approaches have aimed to assist the user by proposing several candidates for this reference frame, and using the demonstration data to decide which candidate is most suitable [3, 4, 5, 6, 7]. However, this still requires defining such candidates beforehand.
A more user-friendly approach would be to derive a suitable task frame directly from the demonstration data itself. To the best of the authors’ knowledge, this has not been explored yet. Most closely related are methods that learn the main directions of motion and force (or compliance) from the data in a given reference frame such as [8, 9, 10, 11, 12]. These main directions could be used to define the orientation of the task frame, as shown in [13] for force directions. However, since the directions are learned in a predefined reference frame, they are fixed to this reference frame, for example, the world or tool frame. Furthermore, deriving an optimal origin for the controller from data, i.e., the origin of the task frame, does not seem to have been considered yet in the literature.
The objective of this paper is to propose a general method to automatically derive an optimal origin and orientation of the task frame from recorded motion and wrench data. To enable non-expert users to more easily learn a large variety of contact tasks, the method needs to have no restriction on the number of degrees of freedom in the task, and no prior assumption on whether the task is dominant in translation or rotation, or whether the task is dominant in force or moment.
@@ I-B Approach
Fig. 1b highlights the step in the LfD workflow that we aim to optimize and automate. To derive the task frame, we first establish two principles for an optimal choice of the task frame based on maximizing the decoupling of signals in the controller. Based on these principles we generate multiple candidates for the origin and the orientation of the task frame from the demonstrated motion and wrench data, assuming that the data are segmented. We also assume that these candidates are either fixed to the world or to the tool. These candidates are processed to obtain the optimal task frame origin and orientation. Once the task frame has been derived, the desired motion and wrench signals are expressed in this frame and a controller can be designed to perform the task.
The automatic task frame derivation procedure is validated for a diverse set of contact-rich tasks, without any assumption on the type of task, by comparing the result with an expert’s choice. The tasks are performed on the robot after designing a constraint-based controller based on the derived task frame. The controller’s robustness and generalization capabilities are tested under disturbances, parameter variations, and alterations in the task frame.
@@ I-C Contribution and Outline
The main contribution of this research is a multi-step algorithm to derive an optimal task frame from demonstrated motion and wrench data. The main discerning features of our approach are that it incorporates data not only of the object’s position and the interaction force, but also of the object’s orientation and the interaction moment. It decides the most appropriate reference frame separately for the origin and orientation of the task. Finally, it decides the most appropriate variable in which the motion and force signals should be parameterized. The method is intended as a tool to automate the selection of a suitable reference frame before learning the task model with some learning method. The intent is to assist users that have limited experience in control.
The remainder of this article is organized as follows. Section II provides a more in-depth review of the literature on contact task learning related to the work proposed in this paper. Section III presents mathematical preliminaries, while Section IV proposes the methodology to automatically derive the optimal task frame from motion and wrench data. Sections V and VI provide an extensive experimental validation of the task frame derivation procedure itself and of the use of the obtained task frame in the performance of learned tasks. Finally, Section VII outlines the major contributions, limitations, and conclusions of the work.
II Relation to the state of the art
@@ II-A Collection of Demonstration Data
Four common ways to collect demonstration data for LfD methods are kinesthetic teaching [11, 14, 15, 13, 8, 3, 7], teleoperation [16, 10], passive observation [17, 18] and simulation. The first method involves physically grasping the robot and moving it through the task, recording motion via the robot’s encoders and interaction wrench via torque sensors at the joints or via a force/torque sensor at the wrist. While cost-effective and fast, this method can be challenging for heavy robots [10], bulky robots [19], redundant robots, dual-arm setups [20], or non-manipulator robots [1], as undesirable dynamics induced by the operator may affect the recorded data [17]. The second method is usually more expensive since it involves an external interface between the robot and the operator, such as a joystick, haptic device or graphical user interface. Collecting demonstration data by teleoperation is also more challenging than kinesthetic teaching [21]. In the third method, the operator uses an instrumented tool to perform the task and record motion and wrench data. In the last method, simulation environments like MoJuCo [22] and Bullet [23] are used to generate demonstration data. The latter method offers abundant data which can enable reinforcement-learning-based methods to widely explore the task without any risk for the robot. However, transferring learned skills from the simulator to the real world is still challenging [24]. Consequently, building a reliable model derived from one [11, 14, 12] or a few [18, 15, 3] demonstrations is a critical aspect of learning contact-rich tasks. In this article, passive observation is employed to record demonstrations as it aligns with the most natural way for the operator to perform tasks [19], but the developed approach is not limited to this demonstration method. Furthermore, the proposed approach to derive an optimal task frame can also be used with simulation data as input or with robot data, i.e., motion and wrench data recorded during a task performed by a robot.
@@ II-B Assumptions and Applications
Some researchers included methods for segmentation in their proposed pipelines for processing the demonstration data, such as [3, 17, 11, 12], but we left segmentation out of the scope of the present paper. I.e., we assume demonstration data has already been segmented and propose an approach to derive a task frame for a single task segment in which the contact situation remains the same over time. On the other hand, while a lot of approaches in literature focus on LfD for a single task or a limited class of tasks, our aim was to develop a generic approach, without further knowledge or assumptions about the task other than that the contact is between only two objects with a measurable relative motion and interaction wrench.
@@ II-C Selection of Task Frame
Task frame selection regained interest in the context of LfD, since demonstration data and task models have to be represented in an appropriate reference frame. There are two common ways to choose a task frame. The first way is, still as in earlier works, to rely on implicit high-level robotics expertise, i.e., without using any prescribed method. The second way is to select between several task frame candidates that are predefined based on expert knowledge. This selection is based on criteria like consistency [3], variance [4] or generalization capability of the demonstration data [6]. A first set of such studies directly select between different task frame candidates [3, 4, 5] or optimize the position and orientation of such candidates [7]. Other studies indirectly learn an appropriate reference frame or components thereof. For example, a method was proposed which learns to activate a set of attractors that are each defined in a different appropriate reference frame [6]. In [13], the authors proposed a method that could adapt the orientation of the task frame w.r.t. the environment. Additionally, some works learned the main directions of motion and force (or compliance) from the data in a given reference frame, such as [8, 9, 10, 11, 12].
In this paper, we propose a generic approach to derive an optimal task frame for contact-rich tasks, allowing interaction with both articulated and non-articulated objects. Both the origin and the orientation of the task frame are derived. The origin and orientation can be fixed in different viewpoints (world or tool) so that the task frame as a whole is not necessarily fixed to either world or tool. We use both motion and wrench data, consider both their translational and rotational vector components and automatically derive the motion and wrench vector components of interest from the demonstration data. The procedure does not involve hyperparameters. While all works mentioned above exhibit some of these features, they lack several others. In particular, to the best of the authors’ knowledge, there is no method available in the literature to derive the task frame from motion and wrench data. Others use application-specific knowledge [30, 8, 9, 17, 16, 31, 10, 32, 13, 18, 33, 15, 11, 12] or predefined candidates [5, 4, 3, 7, 6]. Some works do not consider tasks involving articulated objects [15, 18, 13, 32, 10, 6, 7, 17, 9, 8, 3, 4, 5, 33] or non-articulated objects [12, 11]. Some works do not consider tasks with dominant rotation [3, 8, 9, 17, 16, 31, 6, 10, 32, 13, 15, 11, 12], force [5, 7, 6] or moment [5, 3, 34, 8, 9, 17, 16, 6, 10, 31, 7, 32, 13, 15, 11, 12]. On the other hand, although it is not common to consider a moving task frame for contact-rich tasks in the context of LfD, this was done in [13].
@@ II-D Task Model and Control
The task model used in the experimental validation Sections V and VI is derived from 6-D motion and 6-D wrench data. It consists of 1) a task frame, with its location and orientation as parameters, and 2) the demonstration data, expressed in the task frame and averaged over the demonstration trials to find the desired values for the motion and interaction wrench. No further processing of the data is performed. A generic constraint-based controller is used to follow these desired values. However, we believe that other learning methods could leverage this automatically derived task frame, for building their models and exploring alternative controllers.
@@ III-A Orientation and Pose
Fig. 2 visualizes the similarity transformation under general conditions with three moving frames.
@@ III-B Velocity kinematics
The velocity kinematics of an orientation frame are characterized by the rotational velocity vector $\boldsymbol{\omega}$. When the vector’s coordinates are expressed in $\left[a\right]$, the relation between ${}_{a}\boldsymbol{\omega}$ and the derivative of the rotation matrix $\prescript{b}{}{\vphantom{R}}\dot{\boldsymbol{R}}_{a}$ is:
$\prescript{b}{}{\vphantom{R}}\dot{\boldsymbol{R}}_{a}=[{}_{a}\boldsymbol{\omega}]_{\times}\prescript{b}{}{\boldsymbol{R}}_{a}.$ (10)
The velocity kinematics of a pose frame are characterized by the twist $\boldsymbol{\mathcalboondox{t}}=(\boldsymbol{\omega}^{T}~{}~{}\boldsymbol{v}^{o\,T})^{T}\in\mathbb{R}^{6\times 1}$, consisting of the rotational velocity vector $\boldsymbol{\omega}$ and the translational velocity vector $\boldsymbol{v}^{o}$ at the origin $o$ of the reference frame. When the twist coordinates are expressed in $\{a\}$, the relation between ${}_{a}\boldsymbol{\mathcalboondox{t}}$ and the time-derivative of the pose matrix $\prescript{b}{}{\vphantom{T}}\dot{\boldsymbol{T}}_{a}$ can be worked out as:
$\prescript{b}{}{\vphantom{T}}\dot{\boldsymbol{T}}_{a}=[{}_{a}\boldsymbol{\mathcalboondox{t}}]_{\times}\prescript{b}{}{\boldsymbol{T}}_{a}.$ (11)
@@ III-D Time-invariant motion and wrench trajectories
For a given progress $\xi(t)$ as a function of time, the reparameterizations of pose, wrench, and twist from the time domain to the progress domain are defined as follows [41]:
$\prescript{}{}{\boldsymbol{T}}{}(\xi)=\prescript{}{}{\boldsymbol{T}}{}(t(\xi)),~\boldsymbol{\mathcalboondox{w}}(\xi)={\boldsymbol{\mathcalboondox{w}}}(t(\xi)),~{}\boldsymbol{\mathcalboondox{t}}(\xi)=\frac{\boldsymbol{\mathcalboondox{t}}(t(\xi))}{\dot{\xi}(t(\xi))},$ (15)
where $t(\xi)$ is the inverse function of $\xi(t)$.
We can go back from the progress domain to the time domain using:
$\prescript{}{}{\boldsymbol{T}}{}(t)=\prescript{}{}{\boldsymbol{T}}{}(\xi(t)),~\boldsymbol{\mathcalboondox{w}}(t)={\boldsymbol{\mathcalboondox{w}}}(\xi(t)),~{}\boldsymbol{\mathcalboondox{t}}(t)=\boldsymbol{\mathcalboondox{t}}(\xi(t)){\dot{\xi}(t)}.$ (16)
Consecutive application of (15) and (16) with different progress timings $\xi(t)$ results in a different time trajectory $\prescript{}{}{\boldsymbol{T}}{}(t)$, i.e., the time-dependency of the trajectory has changed, but the spatial path has remained the same.
These re-parameterizations are helpful to master the variability in demonstrations by allowing us to make time-invariant models. In practice, for discrete-time data, these reparameterizations are performed using numerical interpolation. Throughout this paper and in order not to overload notation, we do not explicitly mention the dependence of signals on time or on another progress variable when such dependence is obvious from the context.